OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).

OpenStudy (anonymous):

can you tell me the slope?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

great m=7, x1=1 , y1=-3 this will be your equation for the line: y-y1=m(x-x1)

OpenStudy (anonymous):

plug it in and simplify accordingly

OpenStudy (anonymous):

y - (-3) = 7 (x -1)

OpenStudy (anonymous):

then you subtract -3

OpenStudy (anonymous):

y - (-3) = 7 (x -1) y +3 = 7 (x -1)

OpenStudy (anonymous):

y +3 = 7x - 7 y = 7x - 7-3 y = 7x - 10 yes you do subtract the 3

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

do you do the same for Write the equation of a line parallel to the line y = 3, that passes through the point (1, 6).

OpenStudy (anonymous):

well yes you find the slope.

OpenStudy (anonymous):

and then use the point given to you where x1=1 y1=6 m=slope

OpenStudy (anonymous):

The only thing you need to find is the slope and then you just plug in everything to this equation: y-y1=m(x-x1)

OpenStudy (anonymous):

so it'll b written as this y - 6 = m (x - 1)

OpenStudy (anonymous):

yes all you need is the slope but y=3 has no slope so m=0

OpenStudy (anonymous):

so that means the whole right side becomes zero

OpenStudy (anonymous):

y - 6 = m (x - 1) y - 6 = 0 (x - 1) y - 6 = 0 then you add 6 to both sides y = 6 this is the equation you have now

OpenStudy (anonymous):

oh ohk..yea i worked it out.. i got it now